OpenStudy (anonymous):

log 5 (9th root of 625)?

OpenStudy (anonymous):

log 5 (5^(4/9)) =hence 4/9

OpenStudy (anonymous):

Could you explain in sentences?

OpenStudy (anonymous):

625=5^4 and 9th root of 625 =5^4/9 therefore log 5 (5^(4/9)) = 4/9